# Dual Symmetry Classification of Non-Hermitian Systems and

### v2

@@ Abstract
The study of non-Hermitian systems has gained significant attention in recent years due to its potential for novel physical phenomena, especially relating with topological phase, and practical applications. The eigenenergies of non-Hermitian Hamiltonians generically become complex. For classical systems, non-Hermitian Hamiltonians are frequently employed to describe optical systems with gain and loss[1, 2, 3, 4, 5, 6], topological circuit systems[7, 8, 9, 10, 11, 12, 13], etc. In open quantum systems, a non-Hermitian Hamiltonian can be used to describe the dynamics of a system without quantum jump process[14, 15, 16]. In such cases, the system follows the Schrödinger equation with a non-Hermitian Hamiltonian.
@@
On the other hand, the point gap is unique to non-Hermitian systems and $\mathbb{Z}$ topological invariant defined as the winding number in the complex energy plane for the system with periodic boundary conditions (PBC) predicts that spectra for the system with open boundary conditions (OBC) collapse into arcs within the closed curves of the PBC spectra. The corresponding eigenstates are localized near the open boundaries, known as the non-Hermitian skin effects(NHSE)[15, 30, 31, 32, 33, 34, 35, 34, 25, 36, 37, 38, 39, 40, 41, 42, 43]. One of the famous examples is the Hatano-Nelson model[44, 45, 46]. It is also known that a system with spinful time-reversal symmetry exhibits $\mathbb{Z}_{2}$ topological invariant for the point gap[32].
One promising area of exploration within topology of non-Hermitian physics is the study of non-unitary quantum walks[47, 48, 49, 50]. The discrete-time quantum walk consists of a walker moving on a lattice, obeying a unitary time-evolution operator with each time step[51, 52, 53, 54, 55, 56, 57]. However, the non-unitary quantum walks incorporate gain and loss, resulting in non-Hermitian Hamiltonian dynamics. It is known that the non-unitary time-evolution operator can be classified by translating the symmetry relations of Hamiltonians to time-evolution operators. Some models of non-unitary quantum walks have been realized theoretically[47] and experimentally[48] by temporally alternating photon losses. However, there is no research focused on the topological properties of time-reversal symmetric no-Hermitian system with non-trivial $\mathbb{Z}_{2}$ point-gap topology.
@@
This paper is organized as follows. In Sec.II, we introduce our new way to define symmetry of time-evolution operators in non-unitary quantum walks. In Sec.III, we define a non-unitary quantum walk with spinful time-reversal ${}^{\dagger}$ and discuss the symmetry class of the system. Moreover, we propose methods of experimental realization of our quantum walk. In Sec.IV.1 we calculate the $\mathbb{Z}_{2}$ point gap topological invariant of the non-unitary quantum walk and validate that when the quantum walk possesses a non-trivial $\mathbb{Z}_{2}$ topology, the eigenvalue spectrum of the time evolution operator under reflecting boundary conditions resides within the spectrum under periodic boundary conditions. Simultaneously, the eigenstates are localized near the boundaries of the system, indicating the emergence of the $\mathbb{Z}_{2}$ NHSE. In Sec.IV.2, we study a junction system comprised of two topologically non-trivial quantum walks with inhomogeneous parameters in real space. Intriguingly, we find that both subsystems exhibit skin effects even under periodic boundary conditions.
@@ II Dual symmetry classification of non-Hermitian systems
In Hermitian systems, a system is classified into the 10 AZ symmetry class by considering the presence or absence of three symmetries; time-reversal symmetry (TRS), partical-hole symmetry (PHS) and chiral symmetry (CS) or sublattice symmetry (SLS). In Ref.[17], because of the absence of Hermiticity $(H^{*}\neq H^{T})$, it is shown that the non-Hermitian systems possess the additional symmetry class called $\text{AZ}^{\dagger}$ symmetry class where TRS and PHS are ramified as ${}^{\dagger}$ and ${}^{\dagger}$, respectively. These symmetry relations are summarized as
@@
While it is well known that the classifications based on the above symmetry relations in Eqs. (II) and (II) correctly predicts non-trivial topological phases in non-Hermitian systems, here, we propose alternative way to classify the non-Hermitian systems based on the fact that there is no strict difference in mathematical definition between non-Hermitian Hamiltonians and non-unitary time-evolution operators. By regarding a given non-Hermitian Hamiltonian $H(k)$ as a non-unitary time-evolution operator $U(k)$, the non-Hermitian Hamiltonian can be classified by symmetries defined in Eq. (II). In the same way, regarding a non-unitary time-evolution operator $U(k)$ as an non-Hermitian Hamiltonian $H(k)$, the non-unitary time-evolution operator can be classified by Eq. (II). This means that there are always two ways to classify a given Hamiltonian or a time-evolution operator in non-Hermitian systems, which we call dual symmetry classification. For example, the symmetry class of a give time-evolution operator is determined by the relations in Eqs. (1a)-(1f) or the relations in Eqs. (3a) - (3f). To avoid confusion, we call the former (latter) scheme for the classification as the $H$-type ($U$-type) classification. For the sake of completeness, we summarize all the symmetry relations in Table 1.
Note that Eq. (1b) and Eq. (3a) for ${}^{\dagger}$ are the same. In addition, Eq. (1a) for TRS and Eq. (3d) for ${}^{\dagger}$ are the same. We also note the inverse of the operator on the right-hand side in Eqs. (3a), (3c), and (3f) for $U$-type symmetry requires the strong constraints on the Hamiltonian. However, the minus sign on the right-hand side in Eqs. (1c)-(1f) for $H$-type symmetry requires the time-evolution operator only shifting the quasi-energy by $\pi$.
@@
Symmetry Hamiltonian Time-evolution op. $H$-type $U$-type
@@
${}^{\dagger}$ $\mathcal{T}_{-}H^{T}(k)\mathcal{T}_{-}^{-1}=H(-k)$ $\mathcal{T}_{-}U^{T}(k)\mathcal{T}_{-}^{-1}=U(-k)$ $\mathcal{T}_{-}X^{T}(k)\mathcal{T}_{-}^{-1}=X(-k)\quad\quad$ $\mathcal{T}_{-}X^{T}(k)\mathcal{T}_{-}^{-1}=X(-k)$
@@ III.1 Real space representation
We consider a particle with two internal degrees of freedom, which is called a walker, moving on a pair of parallel chains, $a$ and $b$, which is important to realize ${}^{\dagger}$ with $\mathcal{T}_{-}\mathcal{T}_{-}^{*}=-1$, as we demonstrate later. The state of the walker can be defined as $\left|\psi(t)\right\rangle=\sum_{x,n,s}\psi_{x,n,s}(t)\left|x\right\rangle\otimes\left|n\right\rangle\otimes\left|s\right\rangle$. Here, $|x\rangle$ where $x\in\mathbb{Z}$ represents the position space, $|n=a,b\rangle$ represents the chain space, and $|s=L,R\rangle$ represents the internal state space where two basic vectors of the internal state of the walker is $\left|L\right\rangle=\left(1,0\right)^{T}$ and $\left|R\right\rangle=\left(0,1\right)^{T}$. The dynamics of a quantum walk are described by a time-evolution operator $U$ periodically operating on the wave function of the walker. After $t$ steps, the state of walker can be written as $|\psi(t)\rangle=U^{t}|\psi(0)\rangle$.
@@ III.3 Symmetry
Here, we check the symmetry of the non-unitary time-evolution operator $\tilde{U}(k)$ in Eq. (41) according to our proposal in Sec. II. First, we observe that the operator $\widetilde{U}(k)$ fulfills spinful ${}^{\dagger}$ in $H$ type in Eq. (1b) with $\mathcal{T}_{-}=\tau_{x}\otimes\sigma_{y}$ where $\mathcal{T}_{-}\mathcal{T}_{-}^{*}=-1$. Thus, the operator $\tilde{U}(k)$ also satisfies the ${}^{\dagger}$ in $U$ type in Eq. (3b) since both relations are identical. Due to spinful ${}^{\dagger}$, the eigenvalue of $\tilde{U}(k)$ is always double degenerated.
Furthermore, the operator $\tilde{U}(k)$ satisfies ${}^{\dagger}$ in $H$ type, namely Eq.(1d) with $\Xi_{-}=\tau_{z}\otimes\sigma_{0}$. Thereby, CS in $H$ type in Eq.(1e) with $\Gamma=\tau_{y}\otimes\sigma_{y}$ is also satisfied. Accordingly, the operator $\tilde{U}(k)$ is classified in class ${}^{\dagger}$ in $H$ type, though it is classified in class ${}^{\dagger}$ in $U$ type.
@@
This means that when the operator satisfies Eq. (1d), the operator also satisfies Eq. (42) only for $\text{Re}\left[E_{0}\right]=0$. In other words, ${}^{\dagger}$ is broken for $\text{Im}\left[E_{0}\right]\neq 0$. Accordingly, our quantum walk $\tilde{U}(k)$ belongs in class ${}^{\dagger}$ even in $H$ type if the symmetric point is not a pure imaginary number. While this fact may imply that the dual symmetry classification is not important, this is not the case as we explained in Sec. IV.
@@ IV $\mathbb{Z}_{2}$ point-gap topology of the non-unitary quantum walk
In this section, we study the point-gap topology of our non-unitary quantum walk $\tilde{U}(k)$ in Eq. 41. We recall that the operator $\tilde{U}(k)$ belongs to class ${}^{\dagger}$ in $H$ type for the symmetric point $\text{Re}[E_{0}]=0$, otherwise it belongs to class ${}^{\dagger}$ in $H$ type. On the other hand, the operator $\tilde{U}(k)$ belongs to class ${}^{\dagger}$ in $U$ type for any symmetric point.
@@ IV.1 $\mathbb{Z}_{2}$ point-gap topology under PBC and RBC
The eigenvalues $\lambda$ of $\widetilde{U}(k)$ are shown by red curves in Fig. 3 for two specific values of parameters. Figure 3(a) clearly shows that the point gaps open in finite regions of the symmetric point $E_{0}$, while the point gaps close by slightly changing the value of $\theta_{2}$ from $\pi/5$ to $\pi/6$ as shown in Figure 3(b). However, due to the double degeneracy of the eigenvalue, the energy-winding number for class ${}^{\dagger}$ is always trivial. However, an additional topological invariant, known as $\mathbb{Z}_{2}$ topological invariant, defined in Ref. [17] by
$(-1)^{\nu(\lambda)}=\text{sgn}\left\{\frac{\text{Pf}\left[\left(\widetilde{U}(\pi)-E_{0}\right)\mathcal{T}\right]}{\text{Pf}\left[\left(\widetilde{U}(0)-E_{0}\right)\mathcal{T}\right]}\right\},$ (43)
can be topologically non-trivial for $\text{Re}\left[E_{0}\right]=0$ in class ${}^{\dagger}$ in $H$ type. Here, $\nu\in\{0,1\}$ represents $\mathbb{Z}_{2}$ topological invariant, with $\nu=0$ indicating a topologically trivial phase. We confirmed that $\nu=1$ when $E_{0}$ locates in the closed curves in Fig. 3(a). Here, we emphasize that we use the non-unitary time-evolution operator $\tilde{U}(k)$ instead of the effective Hamiltonian in Eq. (43) to calculate the $\mathbb{Z}_{2}$ topological invariant while Eq. (43) is defined for non-Hermitian Hamiltonians. Therefore, the fact that the regarding the time-evolution operator as the non-Hermitian Hamiltonian, which is crucial for the dual symmetry classification, is essential to study the topological phase.
However, as we mentioned, our quantum walk belongs to class ${}^{\dagger}$ in $H$ type for general value of $E_{0}$. In this case, $\mathbb{Z}_{2}$ topological invariant at a symmetric point $E_{0}$ is defined by
$(-1)^{\nu(\lambda)}$ $=\operatorname{sgn}\left\{\frac{\operatorname{Pf}\left[\left(\widetilde{U}(\pi)-E_{0}\right)\mathcal{T}\right]}{\operatorname{Pf}\left[\left(\widetilde{U}(0)-E_{0}\right)\mathcal{T}\right]}\right.$ (44)
×$\left.\times\exp\left[-\frac{1}{2}\int_{k=0}^{k=\pi}dk\frac{\partial\log\operatorname{det}\left[\left(\widetilde{U}(k)-E_{0}\right)\mathcal{T}\right]}{\partial k}\right]\right\},$
in Ref. [17] Again we remark that we use $\tilde{U}(k)$, not the Hamiltonian, to calculate the $\mathbb{Z}_{2}$ topological invariant. Since Eq. (44) is defined for the operator in $H$ type, if $\mathbb{Z}_{2}$ invariant can be correctly calculated in Eq. (44), this clarifies the correctness of our proposal on the dual symmetry classification.
@@
Here, we check the bulk-edge correspondence for $\mathbb{Z}_{2}$ point-gap topology of our quantum walk. As explained, in case of Fig. 3 (a), the point gap opens in which $\mathbb{Z}_{2}$ topological invariant is nontrivial, while there are no point gap in Fig. 3 (b). To calculate the spectra and eigenstates of the operator $U$ with RBC in the real space representation, we apply for the the generalized Brillouin zone (GBZ) [58, 59, 26, 60] as shown in Fig. 5 (a-1) and (a-2). Figure 3 (a) shows that the spectrum for the system with RBC (RBC spectrum, in short) depicted by blue dots appears as arcs in the point gap. Furthermore, the eigenstates are localized at right and left boundaries, confirming $\mathbb{Z}_{2}$ NHSE. More specifically, as shown in Fig. 5 (a-1) and (a-2), GBZ inside of Brillouin zone (BZ) correspond to eigenmodes localized at the right boundary, while those outside of Brillouin zone (BZ) represent left-boundary-localized eigenmodes [33]. On the other hand, in case of Fig. 3(b) with trivial topology, the RBC spectrum almost overlaps with the PBC spectrum. Moreover, the GBZ coincides with the BZ as shown in Fig. 5(b), resulting in the disappearance of NHSE, and the eigenmodes are not localized near the boundaries as shown in Fig. 5 (d). These numerical results verify that the quantum walk in Eq. (41) exhibits $\mathbb{Z}_{2}$ point-gap topology in $H$ type and satisfies the bulk-edge correspondence for $\mathbb{Z}_{2}$ point-gap topology.

### v4

@@ Abstract
The study of non-Hermitian systems has gained significant attention in recent years due to its potential for novel physical phenomena, especially relating with topological phase, and practical applications[1, 2]. The eigenenergies of non-Hermitian Hamiltonians generically become complex. For classical systems, non-Hermitian Hamiltonians are frequently employed to describe optical systems with gain and loss[3, 4, 5, 6, 7, 8], topological circuit systems[9, 10, 11, 12, 13, 14, 15], etc. In open quantum systems, a non-Hermitian Hamiltonian can be used to describe the dynamics of a system without quantum jump process[16, 17, 18, 19]. In such cases, the system follows the Schrödinger equation with a non-Hermitian Hamiltonian.
@@
On the other hand, the point gap is unique to non-Hermitian systems and $\mathbb{Z}$ topological invariant defined as the winding number in the complex energy plane for the system with periodic boundary conditions (PBC) predicts that spectra for the system with open boundary conditions (OBC) collapse into arcs within the closed curves of the PBC spectra. The corresponding eigenstates are localized near the open boundaries, known as the non-Hermitian skin effects(NHSE)[17, 33, 34, 35, 36, 37, 38, 37, 28, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. One of the famous examples is the Hatano-Nelson model[49, 50, 51]. It is also known that a system with spinful time-reversal symmetry exhibits $\mathbb{Z}_{2}$ topological invariant for the point gap[35].
One promising area of exploration within topology of non-Hermitian physics is the study of non-unitary quantum walks[52, 53, 54, 55, 56, 57]. The discrete-time quantum walk consists of a walker moving on a lattice, obeying a unitary time-evolution operator with each time step[58, 59, 60, 61, 62, 63, 64]. However, the non-unitary quantum walks incorporate gain and loss, resulting in non-Hermitian Hamiltonian dynamics. It is known that the non-unitary time-evolution operator can be classified by translating the symmetry relations of Hamiltonians to time-evolution operators. Some models of non-unitary quantum walks have been considered theoretically[52, 54, 55, 57] and realized experimentally[53, 56] by temporally alternating photon losses. However, there is no research focused on the topological properties of time-reversal symmetric no-Hermitian system with non-trivial $\mathbb{Z}_{2}$ point-gap topology.
@@
This paper is organized as follows. In Sec.II, we introduce our new way to define symmetry of time-evolution operators in non-unitary quantum walks. In Sec.III, we define a non-unitary quantum walk with spinful time-reversal symmetry and discuss the symmetry class of the system. Moreover, we propose methods of experimental realization of our quantum walk. In Sec.IV.1 we calculate the $\mathbb{Z}_{2}$ point gap topological invariant of the non-unitary quantum walk and validate that when the quantum walk possesses a non-trivial $\mathbb{Z}_{2}$ topology, the eigenvalue spectrum of the time evolution operator under reflecting boundary conditions resides within the spectrum under periodic boundary conditions. Simultaneously, the eigenstates are localized near the boundaries of the system, indicating the emergence of the $\mathbb{Z}_{2}$ NHSE. In Sec.IV.2, we study a junction system comprised of two topologically non-trivial quantum walks with inhomogeneous parameters in real space. Intriguingly, we find that both subsystems exhibit skin effects even under periodic boundary conditions.
@@ II Dual symmetry classification of non-Hermitian systems
In Hermitian systems, a system is classified into the 10 AZ symmetry class by considering the presence or absence of three symmetries; time-reversal symmetry (TRS), partical-hole symmetry (PHS) and chiral symmetry (CS) or sublattice symmetry (SLS). In Ref.[20], because of the absence of Hermiticity $(H^{*}\neq H^{T})$, it is shown that the non-Hermitian systems possess the additional symmetry class called $\text{AZ}^{\dagger}$ symmetry class where TRS and PHS are ramified as TRS and PHS, respectively. These symmetry relations are summarized as
@@
While it is well known that the classifications based on the above symmetry relations in Eqs. (1) and (3) correctly predicts non-trivial topological phases in non-Hermitian systems, here, we propose alternative way to classify the non-Hermitian systems based on the fact that there is no strict difference in mathematical definition between non-Hermitian Hamiltonians and non-unitary time-evolution operators. By regarding a given non-Hermitian Hamiltonian $H(k)$ as a non-unitary time-evolution operator $U(k)$, the non-Hermitian Hamiltonian can be classified by symmetries defined in Eq. (3). In the same way, regarding a non-unitary time-evolution operator $U(k)$ as an non-Hermitian Hamiltonian $H(k)$, the non-unitary time-evolution operator can be classified by Eq. (1). This means that there are always two ways to classify a given Hamiltonian or a time-evolution operator in non-Hermitian systems, which we call dual symmetry classification. For example, the symmetry class of a give time-evolution operator is determined by the relations in Eqs. (1a)-(1f) or the relations in Eqs. (3a) - (3f). To avoid confusion, we call the former (latter) scheme for the classification as the $H$-type ($U$-type) classification. For the sake of completeness, we summarize all the symmetry relations in Table 1.
Note that Eq. (1b) and Eq. (3a) for TRS are the same. In addition, Eq. (1a) for TRS and Eq. (3d) for PHS are the same. We also note the inverse of the operator on the right-hand side in Eqs. (3a), (3c), and (3f) for $U$-type symmetry requires the strong constraints on the Hamiltonian. However, the minus sign on the right-hand side in Eqs. (1c)-(1f) for $H$-type symmetry requires the time-evolution operator only shifting the quasi-energy by $\pi$.
@@
Symmetry Hamiltonian Time-evolution op. $H$ type $U$ type
@@
TRS $\mathcal{T}_{-}H^{T}(k)\mathcal{T}_{-}^{-1}=H(-k)$ $\mathcal{T}_{-}U^{T}(k)\mathcal{T}_{-}^{-1}=U(-k)$ $\mathcal{T}_{-}X^{T}(k)\mathcal{T}_{-}^{-1}=X(-k)\quad\quad$ $\mathcal{T}_{-}X^{T}(k)\mathcal{T}_{-}^{-1}=X(-k)$
@@ III.1 Real space representation
We consider a particle with two internal degrees of freedom, which is called a walker, moving on a pair of parallel chains, $a$ and $b$, which is important to realize TRS with $\mathcal{T}_{-}\mathcal{T}_{-}^{*}=-1$, as we demonstrate later. The state of the walker can be defined as $\left|\psi(t)\right\rangle=\sum_{x,n,s}\psi_{x,n,s}(t)\left|x\right\rangle\otimes\left|n\right\rangle\otimes\left|s\right\rangle$. Here, $|x\rangle$ where $x\in\mathbb{Z}$ represents the position space, $|n=a,b\rangle$ represents the chain space, and $|s=L,R\rangle$ represents the internal state space where two basic vectors of the internal state of the walker is $\left|L\right\rangle=\left(1,0\right)^{T}$ and $\left|R\right\rangle=\left(0,1\right)^{T}$. The dynamics of a quantum walk are described by a time-evolution operator $U$ periodically operating on the wave function of the walker. After $t$ steps, the state of walker can be written as $|\psi(t)\rangle=U^{t}|\psi(0)\rangle$.
@@ III.3 Symmetry
Here, we check the symmetry of the non-unitary time-evolution operator $\tilde{U}(k)$ in Eq. (21) according to our proposal in Sec. II. First, we observe that the operator $\widetilde{U}(k)$ fulfills spinful TRS in $H$ type in Eq. (1b) with $\mathcal{T}_{-}=\tau_{x}\otimes\sigma_{y}$ where $\mathcal{T}_{-}\mathcal{T}_{-}^{*}=-1$. Thus, the operator $\tilde{U}(k)$ also satisfies the TRS in $U$ type in Eq. (3b) since both relations are identical. Due to spinful TRS, the eigenvalue of $\tilde{U}(k)$ is always double degenerated.
Furthermore, the operator $\tilde{U}(k)$ satisfies PHS in $H$ type, namely Eq.(1d) with $\Xi_{-}=\tau_{z}\otimes\sigma_{0}$. Thereby, CS in $H$ type in Eq.(1e) with $\Gamma=\tau_{y}\otimes\sigma_{y}$ is also satisfied. Accordingly, the operator $\tilde{U}(k)$ is classified in class DIII in $H$ type, though it is classified in class AII in $U$ type.
@@
This means that when the operator satisfies Eq. (1d), the operator also satisfies Eq. (22) only for $\text{Re}\left[E_{0}\right]=0$. In other words, PHS is broken for $\text{Im}\left[E_{0}\right]\neq 0$. Accordingly, our quantum walk $\tilde{U}(k)$ belongs in class AII even in $H$ type if the symmetric point is not a pure imaginary number. While this fact may imply that the dual symmetry classification is not important, this is not the case as we explained in Sec. IV.
@@ IV $\mathbb{Z}_{2}$ point-gap topology of the non-unitary quantum walk
In this section, we study the point-gap topology of our non-unitary quantum walk $\tilde{U}(k)$ in Eq. 21. We recall that the operator $\tilde{U}(k)$ belongs to class DIII in $H$ type for the symmetric point $\text{Re}[E_{0}]=0$, otherwise it belongs to class AII in $H$ type. On the other hand, the operator $\tilde{U}(k)$ belongs to class AII in $U$ type for any symmetric point.
@@ IV.1 $\mathbb{Z}_{2}$ point-gap topology under PBC and RBC
The eigenvalues $\lambda$ of $\widetilde{U}(k)$ are shown by red curves in Fig. 3 for two specific values of parameters. Figure 3(a) clearly shows that the point gaps open in finite regions of the symmetric point $E_{0}$, while the point gaps close by slightly changing the value of $\theta_{2}$ from $\pi/5$ to $\pi/6$ as shown in Figure 3(b). However, due to the double degeneracy of the eigenvalue, the energy-winding number for class DIII is always trivial. However, an additional topological invariant, known as $\mathbb{Z}_{2}$ topological invariant, defined in Ref. [20] by
$(-1)^{\nu(E_{0})}=\text{sgn}\left\{\frac{\text{Pf}\left[\left(\widetilde{U}(\pi)-E_{0}\right)\mathcal{T}\right]}{\text{Pf}\left[\left(\widetilde{U}(0)-E_{0}\right)\mathcal{T}\right]}\right\},$ (23)
[…]
However, as we mentioned, our quantum walk belongs to class AII in $H$ type for general value of $E_{0}$. In this case, $\mathbb{Z}_{2}$ topological invariant at a symmetric point $E_{0}$ is defined by
$(-1)^{\nu(E_{0})}$ $=\operatorname{sgn}\left\{\frac{\operatorname{Pf}\left[\left(\widetilde{U}(\pi)-E_{0}\right)\mathcal{T}\right]}{\operatorname{Pf}\left[\left(\widetilde{U}(0)-E_{0}\right)\mathcal{T}\right]}\right.$ (24)
×$\left.\times\exp\left[-\frac{1}{2}\int_{k=0}^{k=\pi}dk\frac{\partial\log\operatorname{det}\left[\left(\widetilde{U}(k)-E_{0}\right)\mathcal{T}\right]}{\partial k}\right]\right\}$
[…]
@@
Here, we check the bulk-edge correspondence for $\mathbb{Z}_{2}$ point-gap topology of our quantum walk. As explained, in case of Fig. 3 (a), the point gap opens in which $\mathbb{Z}_{2}$ topological invariant is nontrivial, while there are no point gap in Fig. 3 (b). To calculate the spectra and eigenstates of the operator $U$ with RBC in the real space representation, we apply for the the generalized Brillouin zone (GBZ) [65, 66, 29, 67] as shown in Fig. 5 (a) and (b). Figure 3 (a) shows that the spectrum for the system with RBC (RBC spectrum, in short) depicted by blue dots appears as arcs in the point gap. Furthermore, the eigenstates are localized at right and left boundaries, confirming $\mathbb{Z}_{2}$ NHSE. More specifically, as shown in Fig. 5 (a) and (c), GBZ inside of Brillouin zone (BZ) correspond to eigenmodes localized at the right boundary, while those outside of Brillouin zone (BZ) represent left-boundary-localized eigenmodes [36]. On the other hand, in case of Fig. 3(b) with trivial topology, the RBC spectrum almost overlaps with the PBC spectrum. Moreover, the GBZ coincides with the BZ as shown in Fig. 5(b), resulting in the disappearance of NHSE, and the eigenmodes are not localized near the boundaries as shown in Fig. 5 (d). These numerical results verify that the quantum walk in Eq. (21) exhibits $\mathbb{Z}_{2}$ point-gap topology in $H$ type and satisfies the bulk-edge correspondence for $\mathbb{Z}_{2}$ point-gap topology.